Mina invest her savings of R24 000 in an account that pays 7.5% interest per year, compounded monthly. The amount that she will have in her account after three years is
1
Expert's answer
2020-11-19T10:19:26-0500

"P_{N}= P_{0}(1+\frac{r}{k})^{kN}"

"A_{3}= 24000(1+\frac{0.075}{12})^{3\times 12}"

"=24000(1.00625)^{36} = R 30034.71"